OpenStudy (anonymous):

divide. assume that all variables are nonzero.(65/36) / (25/21) simplify your answer. type as integer or a fraction.

OpenStudy (anonymous):

65 / 36 / 25/21 = 65/36 * 21/25 = 13/12 * 7/5 = 91/60

OpenStudy (anonymous):

or 1 / 31/60